A coin will be tossed twice, and each toss will be recorded as heads (H) or tails (T).

Give the sample space describing all possible outcomes.
Then give all of the outcomes for the event that the second toss is tails.

Use the format

HT

to mean that the first toss is heads and the second is tails.
If there is more than one element in the set, separate them with commas.
Answer #1

Answer:

Here H= Head and T = Tail

1) If a unbiased coin is tossed once only then sample space is \left \{ H, T \right \}

2) If a unbiased coin is tossed twice then sample space is \left \{ HH, HT , TH, TT\right \}

All of the outcomes for the event that the second toss is tails is \left \{ HT , TT\right \}